LjTheBaller16:

The manufacturing cost of a cell phone was $124. Sellers first marked up the cost by 20 percent. Sales were poor, so sellers decreased the markup to 15 percent. Calculate the old and new selling prices.

TETSXPREME:

if the phone costs $ 124 ,and first was marked up by 20 % it means that it was for sale for 124+(124*20)100=124+24.80= $148.80 if the seller decreased the mark up to 15% then the phone costs now 124+(124*15)/100=124+18.60= $ 142.60 ~According To Mr.GooGle~